# Score identity Distillation: Exponentially Fast Distillation of Pretrained Diffusion Models for One-Step Generation

### v1

@@ Abstract
We introduce Score identity Distillation (SiD), an innovative data-free method that distills the generative capabilities of pretrained diffusion models into a single-step generator. SiD not only facilitates an exponentially fast reduction in Fréchet inception distance (FID) during distillation but also approaches or even exceeds the FID performance of the original teacher diffusion models. By reformulating forward diffusion processes as semi-implicit distributions, we leverage three score-related identities to create an innovative loss mechanism. This mechanism achieves rapid FID reduction by training the generator using its own synthesized images, eliminating the need for real data or reverse-diffusion-based generation, all accomplished within significantly shortened generation time. Upon evaluation across four benchmark datasets, the SiD algorithm demonstrates high iteration efficiency during distillation and surpasses competing distillation approaches, whether they are one-step or few-step, data-free, or dependent on training data, in terms of generation quality. This achievement not only redefines the benchmarks for efficiency and effectiveness in diffusion distillation but also in the broader field of diffusion-based generation. Our PyTorch implementation will be publicly accessible on GitHub.
Figure 1: Rapid advancements in the distillation of a pretrained ImageNet 64x64 diffusion model are shown using the proposed SiD method, with settings $\alpha=1.0$, a batch size of 1024, and a learning rate of 5e-6. The series of images, generated from the same set of random noises post-training the SiD generator with varying counts of synthesized images, illustrates progressions at 0, 0.1, 0.2, 0.5, 1, 2, 5, 10, 20, and 50 million images. These are equivalent to roughly 0, 100, 200, 500, 1K, 2K, 5K, 10K, 20K, and 49K training iterations respectively, organized from the top left to the bottom right. The associated FIDs for these iterations are 153.52, 34.83, 37.42, 18.08, 10.82, 7.74, 5.94, 4.49, 3.40, and 3.07, in order. The progression of FIDs is detailed in Fig. 9 in the Appendix.
@@ 2 Related Work
Significant efforts have been directed towards executing the reverse diffusion process in fewer steps. A prominent line of research involves interpreting the diffusion model through the lens of stochastic differential equations (SDE) or ordinary differential equations (ODE), followed by employing advanced numerical solvers for SDE/ODE (Song et al., 2020; Liu et al., 2022a; Lu et al., 2022; Zhang & Chen, 2023; Karras et al., 2022). Despite these advancements, there remains a pronounced trade-off between reducing steps and preserving visual quality. Another line of work considers the diffusion model within the framework of flow matching, applying strategies to simplify the reverse diffusion process into more linear trajectories, thereby facilitating larger step advancements (Liu et al., 2022b; Lipman et al., 2022). To achieve generation within fewer steps, researchers also propose to truncate the diffusion chain and starting the generation from an implicit distribution instead of white Gaussian noise (Pandey et al., 2022; Zheng et al., 2023a; Lyu et al., 2022) and combining it with GANs for faster generation (Xiao et al., 2022; Wang et al., 2023c).
@@ 3 Forward Diffusion as Semi-Implicit Distribution: Exploring Score Identities
The marginal of a mixture distribution can be expressed as $p(\bm{x})=\int p(\bm{x}|\bm{z})p(\bm{z})d\bm{z}.$ In cases where $p(\bm{x}|\bm{z})$ is analytically defined and $p(\bm{z})$ is straightforward to sample from, yet the marginal distribution is intractable or difficult to compute, we follow Yin & Zhou (2018) to refer to it as a semi-implicit distribution. This semi-implicit framework and its derivatives have been widely used to develop flexible variational distributions with tractable parameter inference, as evidenced by a series of studies (Yin & Zhou, 2018; Molchanov et al., 2019; Hasanzadeh et al., 2019; Titsias & Ruiz, 2019; Sobolev & Vetrov, 2019; Lawson et al., 2019; Moens et al., 2021; Yu & Zhang, 2023; Yu et al., 2023).
@@ 3.1 Forward Diffusions and Semi-Implicit Distributions
$\textstyle p_{\text{data}}(\bm{x}_{t})=\int q(\bm{x}_{t}\,|\,\bm{x}_{0})p_{\text{data}}(\bm{x}_{0})\,d\bm{x}_{0},$ (1)
where the forward conditional $q(\bm{x}_{t}\,|\,\bm{x}_{0})$ is analytically defined, but the data distribution $p_{\text{data}}(\bm{x}_{0})$ remains unknown and is typically represented through empirical samples. In this paper, we focus on Gaussian diffusion models, where the forward conditional follows a Gaussian distribution:
$q(\bm{x}_{t}\,|\,\bm{x}_{0})=\mathcal{N}(a_{t}\bm{x}_{0},\sigma_{t}^{2}),$
with $a_{t}\in[0,1]$. To generate a diffused sample from $\bm{x}_{0}\sim p_{\text{data}}(\bm{x}_{0})$, reparameterization is often employed:
$\bm{x}_{t}:=a_{t}\bm{x}_{0}+\sigma_{t}\bm{\epsilon}_{t},\quad\bm{\epsilon}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I}).$
@@
While the exact form of $p_{\text{data}}(\bm{x}_{t})$ and hence the score $S(\bm{x}_{t}):=\nabla_{\bm{x}_{t}}\ln p_{\text{data}}(\bm{x}_{t})$ are not known, the score of the forward conditional $q(\bm{x}_{t}\,|\,\bm{x}_{0})$ has an analytic expression as
$\nabla_{\bm{x}_{t}}\ln q(\bm{x}_{t}\,|\,\bm{x}_{0})=\sigma_{t}^{-2}{(\bm{x}_{0}-\bm{x}_{t})}=-\sigma_{t}^{-1}{\bm{\epsilon}_{t}}.$ (2)
In this work, we explore an implicit generator $p_{\theta}(\bm{x}_{g})$, parameterized by $\theta$, which generates random samples as $\bm{x}_{g}=G_{\theta}(\bm{z}),\bm{z}\sim p(\bm{z}),$ where $G_{\theta}(\cdot)$ represents a neural network, parameterized by $\theta$, that deterministically transforms noise $\bm{z}\sim p(\bm{z})$ into generated data $\bm{x}_{g}$. If the distribution of $p_{\theta}(\bm{x}_{g})$ matches that of $p_{\text{data}}(\bm{x}_{0})$, it then follows that the semi-implicit distribution
$\textstyle p_{\theta}(\bm{x}_{t})=\int q(\bm{x}_{t}\,|\,\bm{x}_{g})p_{\theta}(\bm{x}_{g})\,d\bm{x}_{g}$ (3)
would be identical to $p_{\text{data}}(\bm{x}_{t})$ for any $t$. Conversely, as proved in Wang et al. (2023c), if $p_{\theta}(\bm{x}_{t})$ coincides with $p_{\text{data}}(\bm{x}_{t})$ for any $t$, this implies a match between the generator distribution $p_{\theta}(\bm{x}_{g})$ and the data distribution $p_{\text{data}}(\bm{x}_{0})$.
@@ 3.2 Score Identities
Consider the semi-implicit distribution $p_{\text{data}}(\bm{x}_{t})$ in (1), defined by diffusing real data. The expected value of $\bm{x}_{0}$ given $\bm{x}_{t}$, in line with $q(\bm{x}_{0}\,|\,\bm{x}_{t})=\frac{q(\bm{x}_{t}\,|\,\bm{x}_{0})p_{\emph{\text{data}}}(\bm{x}_{0})}{p_{\emph{\text{data}}}(\bm{x}_{t})}$ as per Bayes’ rule, is connected to the score of $p_{\emph{\text{data}}}(\bm{x}_{t})$ as
$\mathbb{E}[\bm{x}_{0}\,|\,\bm{x}_{t}]=\int\bm{x}_{0}q(\bm{x}_{0}\,|\,\bm{x}_{t})\,d\bm{x}_{0}=\bm{x}_{t}+\sigma_{t}^{2}\nabla_{\bm{x}_{t}}\ln p_{\emph{\text{data}}}(\bm{x}_{t}).$
This identity, known as Tweedie’s Formula (Robbins, 1992; Efron, 2011), provides an estimate for the original data $\bm{x}_{0}$ given $\bm{x}_{t}$, where $\bm{x}_{t}$ is generated as $\bm{x}_{t}\sim\mathcal{N}(\bm{x}_{0},\sigma_{t}^{2}),\bm{x}_{0}\sim p_{\text{data}}(\bm{x}_{0})$. The significance of this relationship has been discussed in Luo (2022) and Chung et al. (2022). An equivalent identity applies to the diffused fake data distribution.
@@
For the semi-implicit distribution $p_{\theta}(\bm{x}_{t})$ defined in (3), resulting from diffusing fake data, the expected value of $\bm{x}_{g}$ given $\bm{x}_{t}$, following $q(\bm{x}_{g}\,|\,\bm{x}_{t})=\frac{q(\bm{x}_{t}\,|\,\bm{x}_{g})p_{\theta}(\bm{x}_{g})}{p_{\theta}(\bm{x}_{t})}$ according to Bayes’ rule, is associated with the score of $p_{\theta}(\bm{x}_{t})$ as
$\mathbb{E}[\bm{x}_{g}\,|\,\bm{x}_{t}]=\int\bm{x}_{g}q(\bm{x}_{g}\,|\,\bm{x}_{t})\,d\bm{x}_{g}=\bm{x}_{t}+\sigma_{t}^{2}\nabla_{\bm{x}_{t}}\ln p_{\theta}(\bm{x}_{t}).$
@@
Given the intractability of $\nabla_{\bm{x}_{t}}\ln p_{\theta}(\bm{x}_{t})$, we introduce a projection vector to estimate the expected value of its product with the score:
$\mathbb{E}_{\bm{x}_{t}\sim p_{\theta}(\bm{x}_{t})}\left[u^{T}(\bm{x}_{t})\nabla_{\bm{x}_{t}}\ln p_{\theta}(\bm{x}_{t})\right]$
$\!\!=\mathbb{E}_{(\bm{x}_{t},\bm{x}_{g})\sim q(\bm{x}_{t}\,|\,\bm{x}_{g})p_{\theta}(\bm{x}_{g})}\left[u^{T}(\bm{x}_{t})\nabla_{\bm{x}_{t}}\ln q(\bm{x}_{t}\,|\,\bm{x}_{g})\right].$
@@
${\mathcal{L}}_{\text{ESM}}=\mathbb{E}_{\bm{x}_{t}\sim p_{\text{data}}(\bm{x}_{t})}\left\|S_{\phi}(\bm{x}_{t})-\nabla_{\bm{x}_{t}}\ln p_{\text{data}}(\bm{x}_{t})\right\|_{2}^{2},$
@@
${\mathcal{L}}_{\text{DSM}}=\mathbb{E}_{q(\bm{x}_{t}\,|\,\bm{x}_{0})p_{\text{data}}(\bm{x}_{0})}\left\|S_{\phi}(\bm{x}_{t})-\nabla_{\bm{x}_{t}}\ln q(\bm{x}_{t}\,|\,\bm{x}_{0})\right\|_{2}^{2}.$
@@ 4.1 Model-based Explicit Score Matching (MESM)
Assuming the existence of a diffusion model for the data, with parameter $\phi$ pretrained to estimate the score $\nabla_{\bm{x}_{t}}\ln p_{\text{data}}(\bm{x}_{t})$, we use the following approximation:
$\textstyle\nabla_{\bm{x}_{t}}\ln p_{\text{data}}(\bm{x}_{t})\approx S_{\phi}(\bm{x}_{t}):={\sigma_{t}^{-2}}({f_{\phi}(\bm{x}_{t},t)-\bm{x}_{t}}).$
In other words, we adopt $f_{\phi}(\bm{x}_{t},t)\approx\mathbb{E}[\bm{x}_{0}|\bm{x}_{t}]$ as our approximation, according to (4) in Identity 1. Our goal is to distill the knowledge encapsulated in $\phi$, extracting which for data generation typically requires many iterations through the same network $f_{\phi}(\bm{\cdot},\bm{\cdot})$.
We use the pretrained score $S_{\phi}(\bm{x}_{t})$ to construct our distillation loss. Our aim is to train $G_{\theta}$ to distill the iterative, multi-step reverse diffusion process into a single network evaluation step. For a specific reverse diffusion time step $t\sim p(t)$, we define the theoretical distillation loss as
${\mathcal{L}}_{\theta}=\mathbb{E}_{\bm{x}_{t}\sim p_{\theta}(\bm{x}_{t})}[\|\delta_{\phi,\theta}(\bm{x}_{t})\|_{2}^{2}],$ (8)
$\delta_{\phi,\theta}(\bm{x}_{t}):=S_{\phi}(\bm{x}_{t})-\nabla_{\bm{x}_{t}}\ln p_{\theta}(\bm{x}_{t}).$ (9)
We refer to $\delta_{\phi,\theta}(\bm{x}_{t})$ as score difference and designate its L2 norm ${\mathcal{L}}_{\theta}$ as the model-based explicit score-matching (MESM) loss, also known in the literature as a Fisher divergence (Lyu, 2009; Holmes & Walker, 2017; Yang et al., 2019; Yu & Zhang, 2023). This differs from the ESM loss defined in (6) as the expectation is computed with respect to the diffused fake data distribution $p_{\theta}(\bm{x}_{t})$ rather than diffused real data distribution $p_{\text{data}}(\bm{x}_{t})$.
A common assumption is that the performance of the student model used for distillation will be limited by the outcomes of reverse diffusion using $S_{\phi}(\bm{x}_{t})$, the teacher model. However, our results demonstrate that the student model, utilizing single-step generation, can indeed exceed the performance of the teacher model, EDM of Karras et al. (2022), which relies on iterative refinement. This indicates that the aforementioned hypothesis might not necessarily hold true. It implies that reverse diffusion could accumulate errors throughout its process, even with very fine-grained reverse steps and the use of advanced numerical solvers designed to counteract error accumulations.

### v2

@@ Abstract
We introduce Score identity Distillation (SiD), an innovative data-free method that distills the generative capabilities of pretrained diffusion models into a single-step generator. SiD not only facilitates an exponentially fast reduction in Fréchet inception distance (FID) during distillation but also approaches or even exceeds the FID performance of the original teacher diffusion models. By reformulating forward diffusion processes as semi-implicit distributions, we leverage three score-related identities to create an innovative loss mechanism. This mechanism achieves rapid FID reduction by training the generator using its own synthesized images, eliminating the need for real data or reverse-diffusion-based generation, all accomplished within significantly shortened generation time. Upon evaluation across four benchmark datasets, the SiD algorithm demonstrates high iteration efficiency during distillation and surpasses competing distillation approaches, whether they are one-step or few-step, data-free, or dependent on training data, in terms of generation quality. This achievement not only redefines the benchmarks for efficiency and effectiveness in diffusion distillation but also in the broader field of diffusion-based generation. The PyTorch implementation is available at https://github.com/mingyuanzhou/SiD.
Diffusion Distillation, Score Matching, Deep Generative Models
Figure 1: Rapid advancements in the distillation of a pretrained ImageNet 64x64 diffusion model are shown using the proposed SiD method, with settings $\alpha=1.0$, a batch size of 1024, and a learning rate of 5e-6. The series of images, generated from the same set of random noises post-training the SiD generator with varying counts of synthesized images, illustrates progressions at 0, 0.1, 0.2, 0.5, 1, 2, 5, 10, 20, and 50 million images. These are equivalent to roughly 0, 100, 200, 500, 1K, 2K, 5K, 10K, 20K, and 49K training iterations respectively, organized from the top left to the bottom right. The associated FIDs for these iterations are 153.52, 34.83, 37.42, 18.08, 10.82, 7.74, 5.94, 4.49, 3.40, and 3.07, in order. The progression of FIDs is detailed in Fig. E in the Appendix.
@@ 2 Related Work
Significant efforts have been directed towards executing the reverse diffusion process in fewer steps. A prominent line of research involves interpreting the diffusion model through the lens of stochastic differential equations (SDE) or ordinary differential equations (ODE), followed by employing advanced numerical solvers for SDE/ODE (Song et al., 2020; Liu et al., 2022a; Lu et al., 2022; Zhang & Chen, 2023; Karras et al., 2022; Xue et al., 2023). Despite these advancements, there remains a pronounced trade-off between reducing steps and preserving visual quality. Another line of work considers the diffusion model within the framework of flow matching, applying strategies to simplify the reverse diffusion process into more linear trajectories, thereby facilitating larger step advancements (Liu et al., 2022b; Lipman et al., 2022). To achieve generation within fewer steps, researchers also propose to truncate the diffusion chain and starting the generation from an implicit distribution instead of white Gaussian noise (Pandey et al., 2022; Zheng et al., 2023a; Lyu et al., 2022) and combining it with GANs for faster generation (Xiao et al., 2022; Wang et al., 2023c).
@@ 3 Forward Diffusion as Semi-Implicit Distribution: Exploring Score Identities
The marginal of a mixture distribution can be expressed as $p(\boldsymbol{x})=\int p(\boldsymbol{x}|\boldsymbol{z})p(\boldsymbol{z})\mathrm{d}\boldsymbol{z}.$ In cases where $p(\boldsymbol{x}|\boldsymbol{z})$ is analytically defined and $p(\boldsymbol{z})$ is straightforward to sample from, yet the marginal distribution is intractable or difficult to compute, we follow Yin & Zhou (2018) to refer to it as a semi-implicit distribution. This semi-implicit framework and its derivatives have been widely used to develop flexible variational distributions with tractable parameter inference, as evidenced by a series of studies (Yin & Zhou, 2018; Molchanov et al., 2019; Hasanzadeh et al., 2019; Titsias & Ruiz, 2019; Sobolev & Vetrov, 2019; Lawson et al., 2019; Moens et al., 2021; Yu & Zhang, 2023; Yu et al., 2023).
@@ 3.1 Forward Diffusions and Semi-Implicit Distributions
$\textstyle p_{\text{data}}(\boldsymbol{x}_{t})=\int q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{0})p_{\text{data}}(\boldsymbol{x}_{0})\,\mathrm{d}\boldsymbol{x}_{0},$ (1)
where the forward conditional $q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{0})$ is analytically defined, but the data distribution $p_{\text{data}}(\boldsymbol{x}_{0})$ remains unknown and is typically represented through empirical samples. In this paper, we focus on Gaussian diffusion models, where the forward conditional follows a Gaussian distribution:
$q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{0})=\mathcal{N}(a_{t}\boldsymbol{x}_{0},\sigma_{t}^{2}\mathbf{I}),$
with $a_{t}\in[0,1]$. To generate a diffused sample from $\boldsymbol{x}_{0}\sim p_{\text{data}}(\boldsymbol{x}_{0})$, reparameterization is often employed:
$\boldsymbol{x}_{t}:=a_{t}\boldsymbol{x}_{0}+\sigma_{t}\boldsymbol{\epsilon}_{t},\quad\boldsymbol{\epsilon}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I}).$
@@
While the exact form of $p_{\text{data}}(\boldsymbol{x}_{t})$ and hence the score $S(\boldsymbol{x}_{t}):=\nabla_{\boldsymbol{x}_{t}}\ln p_{\text{data}}(\boldsymbol{x}_{t})$ are not known, the score of the forward conditional $q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{0})$ has an analytic expression as
$\nabla_{\boldsymbol{x}_{t}}\ln q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{0})=\sigma_{t}^{-2}{(\boldsymbol{x}_{0}-\boldsymbol{x}_{t})}=-\sigma_{t}^{-1}{\boldsymbol{\epsilon}_{t}}.$ (2)
In this work, we explore an implicit generator $p_{\theta}(\boldsymbol{x}_{g})$, parameterized by $\theta$, which generates random samples as $\boldsymbol{x}_{g}=G_{\theta}(\boldsymbol{z}),\boldsymbol{z}\sim p(\boldsymbol{z}),$ where $G_{\theta}(\cdot)$ represents a neural network, parameterized by $\theta$, that deterministically transforms noise $\boldsymbol{z}\sim p(\boldsymbol{z})$ into generated data $\boldsymbol{x}_{g}$. If the distribution of $p_{\theta}(\boldsymbol{x}_{g})$ matches that of $p_{\text{data}}(\boldsymbol{x}_{0})$, it then follows that the semi-implicit distribution
$\textstyle p_{\theta}(\boldsymbol{x}_{t})=\int q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{g})p_{\theta}(\boldsymbol{x}_{g})\,\mathrm{d}\boldsymbol{x}_{g}$ (3)
would be identical to $p_{\text{data}}(\boldsymbol{x}_{t})$ for any $t$. Conversely, as proved in Wang et al. (2023c), if $p_{\theta}(\boldsymbol{x}_{t})$ coincides with $p_{\text{data}}(\boldsymbol{x}_{t})$ for any $t$, this implies a match between the generator distribution $p_{\theta}(\boldsymbol{x}_{g})$ and the data distribution $p_{\text{data}}(\boldsymbol{x}_{0})$.
@@ 3.2 Score Identities
Consider the semi-implicit distribution $p_{\text{data}}(\boldsymbol{x}_{t})$ in (1), defined by diffusing real data. The expected value of $\boldsymbol{x}_{0}$ given $\boldsymbol{x}_{t}$, in line with $q(\boldsymbol{x}_{0}\,|\,\boldsymbol{x}_{t})=\frac{q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{0})p_{\emph{\text{data}}}(\boldsymbol{x}_{0})}{p_{\emph{\text{data}}}(\boldsymbol{x}_{t})}$ as per Bayes’ rule, is connected to the score of $p_{\emph{\text{data}}}(\boldsymbol{x}_{t})$ as
$\mathbb{E}[\boldsymbol{x}_{0}\,|\,\boldsymbol{x}_{t}]=\int\boldsymbol{x}_{0}q(\boldsymbol{x}_{0}\,|\,\boldsymbol{x}_{t})\,\mathrm{d}\boldsymbol{x}_{0}=\boldsymbol{x}_{t}+\sigma_{t}^{2}\nabla_{\boldsymbol{x}_{t}}\ln p_{\emph{\text{data}}}(\boldsymbol{x}_{t}).$
This identity, known as Tweedie’s Formula (Robbins, 1992; Efron, 2011), provides an estimate for the original data $\boldsymbol{x}_{0}$ given $\boldsymbol{x}_{t}$, where $\boldsymbol{x}_{t}$ is generated as $\boldsymbol{x}_{t}\sim\mathcal{N}(\boldsymbol{x}_{0},\sigma_{t}^{2}\mathbf{I}),\boldsymbol{x}_{0}\sim p_{\text{data}}(\boldsymbol{x}_{0})$. The significance of this relationship has been discussed in Luo (2022) and Chung et al. (2022). An equivalent identity applies to the diffused fake data distribution.
@@
For the semi-implicit distribution $p_{\theta}(\boldsymbol{x}_{t})$ defined in (3), resulting from diffusing fake data, the expected value of $\boldsymbol{x}_{g}$ given $\boldsymbol{x}_{t}$, following $q(\boldsymbol{x}_{g}\,|\,\boldsymbol{x}_{t})=\frac{q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{g})p_{\theta}(\boldsymbol{x}_{g})}{p_{\theta}(\boldsymbol{x}_{t})}$ according to Bayes’ rule, is associated with the score of $p_{\theta}(\boldsymbol{x}_{t})$ as
$\mathbb{E}[\boldsymbol{x}_{g}\,|\,\boldsymbol{x}_{t}]=\int\boldsymbol{x}_{g}q(\boldsymbol{x}_{g}\,|\,\boldsymbol{x}_{t})\,\mathrm{d}\boldsymbol{x}_{g}=\boldsymbol{x}_{t}+\sigma_{t}^{2}\nabla_{\boldsymbol{x}_{t}}\ln p_{\theta}(\boldsymbol{x}_{t}).$
@@
Given the intractability of $\nabla_{\boldsymbol{x}_{t}}\ln p_{\theta}(\boldsymbol{x}_{t})$, we introduce a projection vector to estimate the expected value of its product with the score:
$\mathbb{E}_{\boldsymbol{x}_{t}\sim p_{\theta}(\boldsymbol{x}_{t})}\left[u^{T}(\boldsymbol{x}_{t})\nabla_{\boldsymbol{x}_{t}}\ln p_{\theta}(\boldsymbol{x}_{t})\right]$
$\!\!=\mathbb{E}_{(\boldsymbol{x}_{t},\boldsymbol{x}_{g})\sim q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{g})p_{\theta}(\boldsymbol{x}_{g})}\left[u^{T}(\boldsymbol{x}_{t})\nabla_{\boldsymbol{x}_{t}}\ln q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{g})\right].$
@@
${\mathcal{L}}_{\text{ESM}}=\mathbb{E}_{\boldsymbol{x}_{t}\sim p_{\text{data}}(\boldsymbol{x}_{t})}\left\|S_{\phi}(\boldsymbol{x}_{t})-\nabla_{\boldsymbol{x}_{t}}\ln p_{\text{data}}(\boldsymbol{x}_{t})\right\|_{2}^{2},$
@@
${\mathcal{L}}_{\text{DSM}}=\mathbb{E}_{q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{0})p_{\text{data}}(\boldsymbol{x}_{0})}\left\|S_{\phi}(\boldsymbol{x}_{t})-\nabla_{\boldsymbol{x}_{t}}\ln q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{0})\right\|_{2}^{2}.$
@@ 4.1 Model-based Explicit Score Matching (MESM)
Assuming the existence of a diffusion model for the data, with parameter $\phi$ pretrained to estimate the score $\nabla_{\boldsymbol{x}_{t}}\ln p_{\text{data}}(\boldsymbol{x}_{t})$, we use the following approximation:
$\textstyle\nabla_{\boldsymbol{x}_{t}}\ln p_{\text{data}}(\boldsymbol{x}_{t})\approx S_{\phi}(\boldsymbol{x}_{t}):={\sigma_{t}^{-2}}({f_{\phi}(\boldsymbol{x}_{t},t)-\boldsymbol{x}_{t}}).$
In other words, we adopt $f_{\phi}(\boldsymbol{x}_{t},t)\approx\mathbb{E}[\boldsymbol{x}_{0}|\boldsymbol{x}_{t}]$ as our approximation, according to (4) in Identity 1. Our goal is to distill the knowledge encapsulated in $\phi$, extracting which for data generation typically requires many iterations through the same network $f_{\phi}(\boldsymbol{\cdot},\boldsymbol{\cdot})$.
We use the pretrained score $S_{\phi}(\boldsymbol{x}_{t})$ to construct our distillation loss. Our aim is to train $G_{\theta}$ to distill the iterative, multi-step reverse diffusion process into a single network evaluation step. For a specific reverse diffusion time step $t\sim p(t)$, we define the theoretical distillation loss as
${\mathcal{L}}_{\theta}=\mathbb{E}_{\boldsymbol{x}_{t}\sim p_{\theta}(\boldsymbol{x}_{t})}[\|\delta_{\phi,\theta}(\boldsymbol{x}_{t})\|_{2}^{2}],$ (8)
$\delta_{\phi,\theta}(\boldsymbol{x}_{t}):=S_{\phi}(\boldsymbol{x}_{t})-\nabla_{\boldsymbol{x}_{t}}\ln p_{\theta}(\boldsymbol{x}_{t}).$ (9)
We refer to $\delta_{\phi,\theta}(\boldsymbol{x}_{t})$ as score difference and designate its L2 norm ${\mathcal{L}}_{\theta}$ as the model-based explicit score-matching (MESM) loss, also known in the literature as a Fisher divergence (Lyu, 2009; Holmes & Walker, 2017; Yang et al., 2019; Yu & Zhang, 2023). This differs from the ESM loss defined in (6) as the expectation is computed with respect to the diffused fake data distribution $p_{\theta}(\boldsymbol{x}_{t})$ rather than diffused real data distribution $p_{\text{data}}(\boldsymbol{x}_{t})$.
A common assumption is that the performance of the student model used for distillation will be limited by the outcomes of reverse diffusion using $S_{\phi}(\boldsymbol{x}_{t})$, the teacher model. However, our results demonstrate that the student model, utilizing single-step generation, can indeed exceed the performance of the teacher model, EDM of Karras et al. (2022), which relies on iterative refinement. This indicates that the aforementioned hypothesis might not necessarily hold true. It implies that reverse diffusion could accumulate errors throughout its process, even with very fine-grained reverse steps and the use of advanced numerical solvers designed to counteract error accumulations.